solve for y
4y-5/2=3/2y-3/5

Answer by checkley77(12844)

4y-5/2=3/2y-3/5
4y-3y/2=-3/5+5/2
8y/2-3y/2=(-3*2+5*5)/5*2
5y/2=(-6+25)/10
5y/2=19/10
5y*10=2*19
50y=38
y=38/50
y=19/25 ans.
Proof:
4*19/25-5/2=3/2(19/25)-3/5
76/25-5/2=57/50-3/5
(76*2-5*25)/(2*25)=57/50-30/50
(152-125)/50=27/50
27/50=27/50